# MagMax: Leveraging Model Merging for Seamless Continual Learning

### v1

@@ Abstract
This paper introduces a continual learning approach named MagMax, which utilizes model merging to enable large pre-trained models to continuously learn from new data without forgetting previously acquired knowledge. Distinct from traditional continual learning methods that aim to reduce forgetting during task training, MagMax combines sequential fine-tuning with a maximum magnitude weight selection for effective knowledge integration across tasks. Our initial contribution is an extensive examination of model merging techniques, revealing that simple approaches like weight averaging and random weight selection surprisingly hold up well in various continual learning contexts. More importantly, we present MagMax, a novel model-merging strategy that enables continual learning of large pre-trained models for successive tasks. Our thorough evaluation demonstrates the superiority of MagMax in various scenarios, including class- and domain-incremental learning settings.
@@ 2 Related Work
Weights interpolation has recently emerged as an efficient technique for transfer learning that reduces forgetting. After fine-tuning LPM on target data, its weights are interpolated with the weights of (unchanged) LPM, which allows finding a good balance between accuracy on the target domain and zero-shot capabilities of LPM [45]. Such an approach was further extended when merging models across multiple models for OOD performance [44] or in multi-task learning (i.e., Task Vectors [13]). Since then multiple methods have been developed in this area. TIES-Merging [47] reduces the interference when merging models by trimming parameters and electing signs. In [28], the authors linearize the fine-tuning to disentangle weights and facilitate merging. ZipLoRA [35] adapts diffusion models by merging LoRA weights for different styles and subjects. However, those methods were, up-to-date, evaluated on a limited number of scenarios. In this work, we are interested in using those promising approaches to test how they work for different similarities between tasks, as well as when they are compared with simple CL baselines.
@@ 5 Experimental setup
We consider an exemplar-free continual learning scenario in which we cannot store any data from the previous tasks. As a result, we can not tune scaling factor $\lambda$ at merging time as described in [13]. Therefore, we follow no validation scenario from [47] and set constant $\lambda$ for each method based on experiments on CIFAR100/5 setting. We choose $\lambda=0.5$ for MagMax, $\lambda=0.55$ for TIES and $\lambda=1/N$ for Task Vectors. Notice, that choosing $\lambda=1/N$ for Task Vectors simplifies the method to a simple average of task vectors. It makes Task Vectors and Model Soup identical, and we call this method Avg in further experiments. We tune the hyperparameters of CL methods in the same scenario, setting $\lambda=300$ for EWC and $\lambda=0.2$ for LwF.
@@ 6 Main results
Figure 6: Magnitude of task vectors’ parameters are correlated with the consistency of the update direction in the subsequent tasks.
@@
Figure 7: The contribution of parameters is nearly evenly distributed across task when fine-tuning independently. However, for sequential fine-tuning merging prioritizes the later task vectors which accumulated the knowledge about multiple tasks.
@@ 7 Extended analysis
Sequential fine-tuning improves (or at least does not hurt) various merging methods.
@@
Table 6: Starting fine-tuning from the model adapted to a single task ($\theta_{1}$) does not improve the final performance compared to starting from pre-trained weights ($\theta_{0}$). Training sequentially (starting from $\theta_{t-1}$) achieves the best results. We report the accuracy (%) after the final task and bold the best results.
@@ Acknowledgments
Daniel Marczak is supported by National Centre of Science (NCN, Poland) Grant No. 2021/43/O/ST6/02482. This research was partially funded by National Science Centre, Poland, grant no 2020/39/B/ST6/01511 and grant no 2022/45/B/ST6/02817. Bartłomiej Twardowski acknowledges the grant RYC2021-032765-I. This paper has been supported by the Horizon Europe Programme (HORIZON-CL4-2022-HUMAN-02) under the project "ELIAS: European Lighthouse of AI for Sustainability", GA no. 101120237. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016393.
Appendix 0.A Additional results
@@
Appendix 0.B Distribution of parameters in task vectors

### v2

@@ Abstract
This paper introduces a continual learning approach named MagMax, which utilizes model merging to enable large pre-trained models to continuously learn from new data without forgetting previously acquired knowledge. Distinct from traditional continual learning methods that aim to reduce forgetting during task training, MagMax combines sequential fine-tuning with a maximum magnitude weight selection for effective knowledge integration across tasks. Our initial contribution is an extensive examination of model merging techniques, revealing that simple approaches like weight averaging and random weight selection surprisingly hold up well in various continual learning contexts. More importantly, we present MagMax, a novel model-merging strategy that enables continual learning of large pre-trained models for successive tasks. Our thorough evaluation demonstrates the superiority of MagMax in various scenarios, including class- and domain-incremental learning settings. The code is available at https://github.com/danielm1405/magmax.
@@ 2 Related Work
Weights interpolation has recently emerged as an efficient technique for transfer learning that reduces forgetting. After fine-tuning LPM on target data, its weights are interpolated with the weights of (unchanged) LPM, which allows finding a good balance between accuracy on the target domain and zero-shot capabilities of LPM [46]. Such an approach was further extended when merging models across multiple models for OOD performance [45] or in multi-task learning (i.e., Task Vectors [13]). Since then multiple methods have been developed in this area. TIES-Merging [48] reduces the interference when merging models by trimming parameters and electing signs. In [29], the authors linearize the fine-tuning to disentangle weights and facilitate merging. ZipLoRA [36] adapts diffusion models by merging LoRA weights for different styles and subjects. However, those methods were, up-to-date, evaluated on a limited number of scenarios. In this work, we are interested in using those promising approaches to test how they work for different similarities between tasks, as well as when they are compared with simple CL baselines. A concurrent work, CoFiMA [24], utilizes Fisher Merging [26] sequentially after each task to continually train closed vocabulary image classifiers. In contrast, we focus on reducing parameter-level interferences in open vocabulaty models.
@@ 5 Experimental setup
We consider an exemplar-free continual learning scenario in which we cannot store any data from the previous tasks. As a result, we can not tune scaling factor $\lambda$ at merging time as described in [13]. Therefore, we follow no validation scenario from [48] and set constant $\lambda$ for each method based on experiments on CIFAR100/5 setting. We choose $\lambda=0.5$ for MagMax, $\lambda=0.55$ for TIES and $\lambda=1/N$ for Task Vectors. Notice, that choosing $\lambda=1/N$ for Task Vectors simplifies the method to a simple average of task vectors. It makes Task Vectors and Model Soup identical, and we call this method Avg in further experiments. We tune the hyperparameters of CL methods in the same scenario, setting $\lambda=1e6$ for EWC and $\lambda=0.3$ for LwF.
@@ 6 Main results
Figure 6: Magnitude of parameters of sequentially fine-tuned task vectors is correlated with the consistency of the update direction in the subsequent tasks. We report the results in CIFAR100/10 setting.
Figure 7: The contribution of parameters is nearly evenly distributed across task when fine-tuning independently. However, for sequential fine-tuning merging prioritizes the later task vectors which accumulated the knowledge about multiple tasks. We report the results in CIFAR100/10 setting.
@@ 7 Extended analysis
Sequential fine-tuning improves various merging methods.
@@
Table 6: Starting fine-tuning from the model adapted to a single task ($\theta_{1}$) does not improve the final performance compared to starting from pre-trained weights ($\theta_{0}$). Training sequentially (starting from $\theta_{t-1}$) achieves the best results.
@@ Acknowledgments
Daniel Marczak is supported by National Centre of Science (NCN, Poland) Grant No. 2021/43/O/ST6/02482. This research was partially funded by National Science Centre, Poland, grant no: 2020/39/B/ST6/01511, 2022/45/B/ST6/02817 and 2023/51/D/ST6/02846. Bartłomiej Twardowski acknowledges the grant RYC2021-032765-I. This paper has been supported by the Horizon Europe Programme (HORIZON-CL4-2022-HUMAN-02) under the project "ELIAS: European Lighthouse of AI for Sustainability", GA no. 101120237. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016393.
Appendix 0.A More results
0.A.1 CIL results with different backbones
@@
0.A.2 Sign conflicts
Fig. 10 presents the sign conflicts for class-incremental, domain-incremental and 8 datasets scenarios. We observe that sequential fine-tuning significantly reduces sign conflicts similarly to CIL results presented in Figure 3 in the main paper.
Figure 10: Sign conflicts for CIL, DIL and 8 datasets settings.
0.A.3 Task agnostic per-task results
Figure 11 presents more per-task task-agnostic results with MagMax.
Figure 11: Task-agnostic results of MagMax in different settings.
Appendix 0.B Additional analyses
0.B.1 Layer-wise weight changes
To better undestand the process of fine-tuning and merging with MagMax, we analyze the magnitudes of $\tau_{\textsc{MagMax}_{t}}$ parameters. We group these parameters either by their type (e.g. layer normalization, attention or MLP) or by the block index to which they belong. We present the analysis in Figure 12 and observe that the magnitudes of layer normalization are much higher that the magnitudes of other layers. Moreover, magnitude seems not to depend on the depth. Note, that we only analyze weight matrices and disregard the biases.
Figure 12: Mean magnitudes of $\tau_{\textsc{MagMax}_{t}}$ parameters grouped by layer type (top) or block index (bottom) for CIFAR100/5 (left) and ImageNetR/10 (right). Top: parameters of LayerNorm layers change the most. Bottom: the magnitude of parameter change does not depend much on a block index (depth).
0.B.2 Distribution of parameters in task vectors